# General Preference Modeling with Preference Representations for Aligning Language Models

### v1

@@ 7 Conclusion
This work introduced preference representation learning, a framework for modeling human preferences that can capture complex, intransitive structures like cyclic preferences. The proposed General Preference representation model (GPM) achieves linear complexity while maintaining the ability to model intricate preference relationships. It consistently outperforms traditional models like Bradley-Terry and supervised preference models across various benchmarks, including cyclic preference datasets and real-world tasks from RewardBench. Additionally, incorporating preference scores from GPM into policy optimization methods, such as SPPO and the newly introduced General Preference Optimization (GPO), led to significant performance improvements in downstream tasks that require alignment with intricate human preferences, as demonstrated in benchmarks like AlpacaEval 2.0 and MT-Bench.
Limitations.
Despite the current interesting experimental results, further exploration is still needed. The choice of representation (embedding) dimensions and model architecture can influence the expressiveness and computational efficiency of GPM. An overly expressive model such as a high embedding dimension can impact performance, particularly in tasks prone to overfitting. Future works should examine the expressiveness of preference representations in a wider range of real-world datasets.
Appendix A More on General Preference Representation Learning
In this section, we present additional discussion on general preference modeling with preference representations.
@@ Appendix A More on General Preference Representation Learning
A.1 Complex Representations Interpretation
Our model can also be interpreted using complex representations. By representing the representations as complex vectors $\mathbf{v}_{\mathbf{y}}\in\mathbb{C}^{k}$, we can express the preference score as:
@@ A.1 Complex Representations Interpretation
Theorem A.2 (Expressiveness of Complex Preference Representations).
@@
A.2 Implementing General Preference Representation Model
Theorem A.3 (Expressiveness of Preference Representation Model).
Let $\mathbf{P}\in\mathbb{R}^{2k\times 2k}$ be a real skew-symmetric matrix (i.e., $\mathbf{P}=-\mathbf{P}^{\top}$). Then there exist representations (embeddings) $\{\mathbf{v}_{i}\}_{i=1}^{2k}\subset\mathbb{R}^{2k}$ and a block-diagonal skew-symmetric matrix $\mathbf{R}^{\succ}\in\mathbb{R}^{2k\times 2k}$, with $\mathbf{R}^{\succ}$ consisting of $k$ blocks of the form:
$\mathbf{R}_{l}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix},\quad l=1,\dots,k,$
such that:
To effectively capture general preferences while maintaining computational efficiency, we implement our preference representation model by augmenting an existing language model with two additional components: an eigenvalue scale gate and an eigenvector embedding head.
$\{\lambda_{l}(\mathbf{x})\}=\mathcal{G}_{\lambda}(\mathbf{x}).$
where $\mathbf{D}(\mathbf{x})$ is a block-diagonal matrix with blocks $\sqrt{\lambda_{l}(\mathbf{x})}\mathbf{I}_{2}$, and $\mathbf{R}^{\succ}$ is the skew-symmetric operator defined previously. We normalize the embeddings $\mathbf{v}_{\mathbf{y}}$ to have unit length to ensure stability in training.
This implementation provides a scalable and flexible way to model complex human preferences, making it suitable for large-scale applications in language modeling and other domains where alignment with nuanced human values is essential.
A.3 Training Objective
@@
$\mathcal{L}_{\text{CE}}=-\sum_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\in\mathcal{D}}$ $\left[\mathbb{P}_{\mathcal{D}}(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})\log\sigma\left(\frac{1}{\beta}s(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})\right)\right.$
@@
A.4 Appendix for Proofs
Proof of the Proposition A.1.
@@ 2. Magnitude preserving:
$\mathbf{a}_{i}^{\top}\mathbf{b}_{j}$ $=\mathbf{e}_{i}^{\top}\mathbf{b}_{j}=(\mathbf{b}_{j})_{i}=\frac{1}{2}P_{ji}=-\frac{1}{2}P_{ij},$ (A.1)
$\mathbf{b}_{i}^{\top}\mathbf{a}_{j}$ $=\mathbf{b}_{i}^{\top}\mathbf{e}_{j}=(\mathbf{b}_{i})_{j}=\frac{1}{2}P_{ij}.$ (A.2)
@@
Proof of the Theorem A.2.
@@
Proof of the Theorem A.3.
Proof.
$\mathbf{P}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$
$\mathbf{J}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix}.$
Observe that $\bm{\Lambda}=\mathbf{D}\mathbf{R}^{\succ}\mathbf{D}$.
Set $\mathbf{V}=\mathbf{U}\mathbf{D}$. Then,
Therefore,
where $\mathbf{v}_{i}$ is the $i$-th row of $\mathbf{V}$.
Appendix B More on Experiments
B.1 Additional Experimental Results
Table 7: Impact of the embedding head and the scale gate on the GPM’s performance on RewardBench. Dim. represents the dimension of the embedding head. The highest average scores for each base model are in bold and the second highest are underlined.
@@ B.1 Additional Experimental Results
w. scale gate w. l2 2 68.4 66.7 70.7 80.3 71.5
w. scale gate w.o. l2 2 70.7 66.4 72.8 71.4 70.3
w. o. scale gate w. l2 2 68.4 62.9 70.3 72.5 68.5
w. o. scale gate w.o. l2 2 67.0 69.5 70.2 68.1 68.7
w. scale gate w. l2 4 63.1 68.4 72.4 81.0 71.2
w. scale gate w.o. l2 4 69.3 71.7 67.1 73.8 70.5
w. o. scale gate w. l2 4 71.2 60.7 66.6 76.6 68.8
w. o. scale gate w.o. l2 4 66.8 67.8 71.7 69.1 68.9
w. scale gate w. l2 6 68.2 66.4 80.6 76.7 73.0
w. scale gate w.o. l2 6 66.8 62.9 68.5 62.8 65.2
w. o. scale gate w. l2 6 65.4 65.1 66.9 63.6 65.3
w. o. scale gate w.o. l2 6 69.8 71.1 73.3 74.9 72.3
w. scale gate w. l2 8 71.5 69.7 81.3 75.6 74.5
w. scale gate w.o. l2 8 69.8 66.9 78.9 75.9 72.9
w. o. scale gate w. l2 8 67.3 66.2 62.9 70.8 66.8
w. o. scale gate w.o. l2 8 69.6 67.8 76.1 70.4 71.0
@@
w. scale gate w. l2 2 92.7 87.3 90.5 96.1 91.7
w. scale gate w.o. l2 2 92.5 87.1 88.4 96.1 91.0
w. o. scale gate w. l2 2 92.2 88.6 90.6 97.4 92.2
w. o. scale gate w.o. l2 2 93.0 87.1 88.4 94.3 90.7
w. scale gate w. l2 4 93.6 87.1 90.4 96.0 91.8
w. scale gate w.o. l2 4 93.0 87.9 91.1 96.1 92.0
w. o. scale gate w. l2 4 93.9 87.3 90.8 96.7 92.2
w. o. scale gate w.o. l2 4 93.6 87.5 88.4 94.5 91.0
w. scale gate w. l2 6 93.3 88.6 90.6 96.0 92.1
w. scale gate w.o. l2 6 93.9 87.9 89.3 95.7 91.7
w. o. scale gate w. l2 6 93.9 85.5 91.5 96.3 91.8
w. o. scale gate w.o. l2 6 94.4 89.3 87.6 93.4 91.2
w. scale gate w. l2 8 92.7 86.4 90.4 96.7 91.6
w. scale gate w.o. l2 8 94.1 87.3 89.7 95.5 91.7
w. o. scale gate w. l2 8 93.0 87.1 91.1 96.6 92.0
w. o. scale gate w.o. l2 8 93.9 88.2 89.7 94.9 91.7
Ablations on Scale Gate and Embedding head. We investigate the effects of scale gates and embedding head dimensions, with and without L2 normalization, on model performance. As shown in Table 7 , for Gemma-2B-it models, incorporating a scale gate generally enhances GPM performance across various embedding dimensions. L2 normalization on the embedding head output consistently improves models with scale gates. Interestingly, Gemma-2B-it-based models without L2 normalization or scale gates outperform those with L2 normalization but no scale gates. A plausible explanation for this phenomenon is that removing L2 normalization introduces additional degrees of freedom, particularly beneficial for models with smaller parameter spaces and high-dimensional embedding layers. This increased flexibility may allow the model to better utilize its limited parametric capacity, potentially leading to enhanced expressiveness and task-specific adaptability.
For larger models, such as those based on Llama3.1-8B-Instruct, the impact of scale gates becomes less pronounced. This diminished effect may be attributed to the inherently stronger representational capacity of the 8B parameter model, which can likely capture complex patterns more effectively without additional architectural modifications.
These observations suggest a nuanced relationship between model size, normalization techniques, and architectural enhancements like scale gates, highlighting the importance of considering these factors in model design and optimization.
More Results on Evaluating Language Model Alignment. We further conduct a rigorous evaluation of our downstream task-specific models using various benchmarks. AlpacaEval 2.0 evaluation results are listed in Table 4 and Table 5, using GPT-4-turbo and Deepseek-V2 as evaluators respectively. For LM-Harness, we chose Arc-Challenge, TruthfulQA, WinoGrande, GSM8k, HellaSwag, and MMLU as the evaluation tasks, and used the default rule-based evaluator of lm-evaluation-harness for accuracy calculation. These tasks are the same as those evaluated by Open LLM Leaderboard v1 (Beeching et al., 2023), which no longer provides service. Notice that the evaluator of LM-Harness is not the same as Open LLM Leaderboard v1, so we only compare different models within our results and do not compare them with the Open LLM Leaderboard v1. The results are showed in Tables 8 and 9. To facilitate direct comparison with current state-of-the-art models, we adhere to the evaluation protocol established by the Open LLM Leaderboard v1. Our models are evaluated locally using this standardized framework. The resultant performance metrics are presented in Tables 11 and Table 10.
Table 8: LM-Harness evaluation results of LLama3-8B-it model fine-tuned using SPPO with BT reward model and our GPM.
Size Type Iter SPPO
Arc TruthfulQA WinoGrande GSM8k HellaSwag MMLU Average
base 56.91 51.65 71.43 32.83 75.78 63.72 58.72
2B BT RM 1 55.89 52.94 71.82 26.46 75.95 63.79 57.81
2 54.69 52.61 71.74 21.46 75.69 63.89 56.68
3 53.41 52.23 71.82 19.56 75.70 63.91 56.10
GPM 1 55.20 52.92 71.59 26.99 75.64 63.84 57.70
2 53.84 53.12 71.51 23.88 75.37 63.65 56.90
3 53.50 52.98 71.82 28.89 75.22 63.47 57.65
8B BT RM 1 56.14 55.18 71.51 25.93 76.35 63.99 58.18
2 56.23 56.32 72.22 27.45 76.19 63.77 58.70
3 56.06 55.80 72.61 25.78 76.08 63.87 58.37
GPM 1 55.89 55.40 71.67 28.35 76.35 63.89 58.59
2 55.12 55.53 71.90 33.13 76.01 63.84 59.25
3 55.12 55.66 72.06 29.87 75.89 64.05 58.78

### v2

@@ 7 Conclusion
In this work, we introduce preference embedding, a framework for modeling human preferences that can capture complex, intransitive structures. Our General Preference embedding model (GPM) achieves linear complexity while can model intricate preference relationships. It consistently outperforms traditional models like Bradley-Terry reward models across various benchmarks, including cyclic preference datasets and real-world tasks from RewardBench. Additionally, incorporating preference scores from GPM into policy optimization methods, such as SPPO and the newly introduced General Preference Optimization (GPO), led to performance improvements in downstream tasks that require alignment with intricate human preferences.
Appendix A More on General Preference Embedding
In this section, we present additional discussion on general preference modeling with preference embeddings.
@@ Appendix A More on General Preference Embedding
A.1 Complex Embeddings Interpretation
Our model can also be interpreted using complex embeddings. By representing the embeddings as complex vectors $\mathbf{v}_{\mathbf{y}}\in\mathbb{C}^{k}$, we can express the preference score as:
@@ A.1 Complex Embeddings Interpretation
Theorem A.2 (Expressiveness of Complex Preference Embeddings).
@@
A.2 Training Objective
@@
$\mathcal{L}_{\text{CE}}=-\sum_{(\mathbf{x},\mathbf{y}_{w},\mathbf{y}_{l})\in\mathcal{D}}$ $\left[\mathbb{P}_{\mathcal{D}}(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})\log\sigma\left(\frac{1}{\beta}s(\mathbf{y}_{w}\succ\mathbf{y}_{l}\mid\mathbf{x})\right)\right.$ (A.1)
@@
Appendix B More on General Preference Optimization
Note that General Preference Optimization (GPO) employs an iterative framework inspired by the multiplicative weights update (MWU) algorithm (Freund & Schapire, 1999), which update rule is formulated as:
$\pi_{t+1}(\mathbf{y}\mid\mathbf{x})\propto\pi_{t}(\mathbf{y}\mid\mathbf{x})\exp\left(\eta\cdot s\left(\mathbf{y}\succ\pi_{t}\mid\mathbf{x}\right)\right),$
$\quad t=1,2,\dots,$
where $\eta$ denotes the learning rate and $s\left(\mathbf{y}\succ\pi_{t}\mid\mathbf{x}\right)$ represents the preference score of response $\mathbf{y}$ over the current policy $\pi_{t}$ given prompt $\mathbf{x}$.
The von Neumann winner represents a fundamental concept in social choice theory (Sen, 1986) that has found significant applications in preference-based reinforcement learning (Owen, 2013; Dudík et al., 2015). It corresponds to the Nash equilibrium of a two-player symmetric game (3.2), representing a mixed strategy—a probability distribution over possible responses—that performs optimally against any opponent in the worst-case scenario.
For notational clarity, we define the preference score of a policy $\pi$ over another policy $\pi^{\prime}$ as:
$s\left(\pi\succ\pi^{\prime}\mid\mathbf{x}\right)=\mathbb{E}_{\mathbf{y}\sim\pi(\cdot\mid\mathbf{x}),\ \mathbf{y}^{\prime}\sim\pi^{\prime}(\cdot\mid\mathbf{x})}\left[s\left(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x}\right)\right].$
A distribution $\pi^{*}$ is formally defined as a von Neumann winner when it satisfies:
$\min_{\pi^{\prime}\in\Delta}\ \mathbb{E}_{\mathbf{x}\sim\mathcal{X}}\left[s\left(\pi^{*}\succ\pi^{\prime}\mid\mathbf{x}\right)\right]\geq 0.$
This condition ensures that $\pi^{*}$ is, on average, at least as preferred as any other policy $\pi^{\prime}$. The symmetric nature of the two-player game (3.2) guarantees the existence of such a winner.
Appendix C Proofs of Theorems
C.1 Proof of Proposition A.1
@@ 2. Magnitude preserving:
$\mathbf{a}_{i}^{\top}\mathbf{b}_{j}$ $=\mathbf{e}_{i}^{\top}\mathbf{b}_{j}=(\mathbf{b}_{j})_{i}=\frac{1}{2}P_{ji}=-\frac{1}{2}P_{ij},$ (C.1)
$\mathbf{b}_{i}^{\top}\mathbf{a}_{j}$ $=\mathbf{b}_{i}^{\top}\mathbf{e}_{j}=(\mathbf{b}_{i})_{j}=\frac{1}{2}P_{ij}.$ (C.2)
@@
C.3 Proof of Theorem 4.5.
Proof.
$\mathbf{P}=\mathbf{U}\bm{\Lambda}\mathbf{U}^{\top},$
$\mathbf{J}=\begin{bmatrix}0&-1\\ 1&0\\ \end{bmatrix}.$
Observe that $\bm{\Lambda}=\mathbf{D}\mathbf{R}^{\succ}\mathbf{D}$.
Set $\mathbf{V}=\mathbf{U}\mathbf{D}$. Then,
Therefore,
where $\mathbf{v}_{i}$ is the $i$-th row of $\mathbf{V}$.
C.4 Proof of Theorem 5.1
Proof.
First, since the preference score $s$ is bounded in $[-\rho,\rho]$, we can normalize it to $[0,1]$ by the transformation:
$\widetilde{s}(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x})=\frac{s(\mathbf{y}\succ\mathbf{y}^{\prime}\mid\mathbf{x})}{2\rho}+\frac{1}{2}$
By Theorem 1 in Freund & Schapire (1999), for any sequence of mixed policies $\mu_{1},\mu_{2},\ldots,\mu_{T}$, the sequence of policies $\pi_{1},\pi_{2},\ldots,\pi_{T}$ produced by GPO satisfies:
$\sum_{t=1}^{T}\widetilde{s}(\pi_{t}\prec\mu_{t})\leq\min_{\pi}\left[\frac{\eta}{1-e^{-\eta}}\sum_{t=1}^{T}\widetilde{s}(\pi\prec\mu_{t})+\frac{\mathrm{KL}(\pi\|\pi_{0})}{1-e^{-\eta}}\right]$
Setting $\mu_{t}=\pi_{t}$, note that $\widetilde{s}(\pi_{t}\prec\pi_{t})=\frac{1}{2}$ due to the normalization and symmetry. Thus:
$\frac{T}{2}\leq\min_{\pi}\left[\frac{\eta T}{1-e^{-\eta}}\widetilde{s}(\pi\prec\bar{\pi}_{T})+\frac{\mathrm{KL}(\pi\|\pi_{0})}{1-e^{-\eta}}\right]$
where $\bar{\pi}_{T}=\frac{1}{T}\sum_{t=1}^{T}\pi_{t}$ is the mixture policy.
Rearranging terms:
$\frac{1-e^{-\eta}}{2\eta}\leq\min_{\pi}\left[\widetilde{s}(\pi\prec\bar{\pi}_{T})+\frac{\mathrm{KL}(\pi\|\pi_{0})}{\eta T}\right]$
Since $\pi_{0}$ is an autoregressive model with finite vocabulary support, $|\log\pi_{0}(\cdot)|$ is bounded from above. Thus:
$\mathrm{KL}(\pi\|\pi_{0})\leq\|\log\pi_{0}(\cdot)\|_{\infty}$
Setting $\eta=\frac{\|\log\pi_{0}(\cdot)\|_{\infty}}{\sqrt{T}}$ and using Taylor expansion $\frac{1-e^{-\eta}}{2\eta}=\frac{1}{2}-\frac{\eta}{4}+O(\eta^{2})$:
$\frac{1}{2}-\frac{\|\log\pi_{0}(\cdot)\|_{\infty}}{4\sqrt{T}}+O(T^{-1})\leq\min_{\pi}\left[\widetilde{s}(\pi\prec\bar{\pi}_{T})\right]+\sqrt{\frac{\|\log\pi_{0}(\cdot)\|_{\infty}}{T}}$
Converting back to the original preference score scale:
$\min_{\pi}\left[s(\pi\prec\bar{\pi}_{T})\right]\geq-\frac{\rho}{2}-O\left(\frac{\rho}{\sqrt{T}}\right)$
By symmetry:
$\max_{\pi}\left[s(\pi\succ\bar{\pi}_{T})\right]\leq\frac{\rho}{2}+O\left(\frac{\rho}{\sqrt{T}}\right)$
Therefore, the duality gap is:
$\max_{\pi}s(\pi\succ\bar{\pi}_{T})-\min_{\pi}s(\pi\prec\bar{\pi}_{T})$
$=\max_{\pi}s(\pi\succ\bar{\pi}_{T})-\min_{\pi}s(\pi\prec\bar{\pi}_{T})$
$=O\left(\frac{1}{\sqrt{T}}\right)$
Proof of Theorem A.2
@@
Appendix D More Related Work
Intransitivity in Game Theory. The symmetric zero-sum game and its intransitivity have also been frequently studied in the context of game theory. Balduzzi et al. (2018) was motivated by evaluation among different agents, showing that any symmetric zero-sum game can be decomposed into a “transitive” game and a “cyclic” game, and proposed Nash averaging for better agent/task evaluation. Balduzzi et al. (2019) generalized the results from matrix games to functional-form games and propose new algorithms to construct diverse populations of effective agents. Czarnecki et al. (2020) investigated the geometrical properties of real-world games (e.g., Tic-Tac-Toe, Go, StarCraft II) and proposed that real-world games have a “spinning top” geometry, with a strong transitive dimension and gradually diminishing non-transitive cyclic dimensions. Very recently, Bertrand et al. (2023) examined the limitations of the Elo rating system and proposed an alternative “disc decomposition” method that can better handle both transitive and cyclic game dynamics.
Representation Learning and Embedding. Representation learning and embedding techniques have successfully captured relational structures across various domains (Mikolov et al., 2013; Chen et al., 2020; Radford et al., 2021), yet their application in preference modeling and RLHF remains limited. Our work introduces preference representation learning, an approach that enhances expressiveness while maintaining computational efficiency, bridging the gap left by traditional approaches.
Appendix E More on Experiments
Cyclic Preference Dataset. We constructed a dataset by inducing cyclic preferences from the Ultrafeedback dataset Cui et al. (2024). The dataset includes responses evaluated across four key metrics: instruction following, honesty, truthfulness, and helpfulness. We created preference cycles such as: instruction following $\succ$ honesty $\succ$ truthfulness $\succ$ helpfulness $\succ$ instruction following, ensuring the presence of intransitive cycles. We further generated four sub-datasets by omitting one metric from each cycle, resulting in 4 different datasets with 216 to 363 instances.
E.1 Additional Ablation Studies
Ablations on Scale Gate and Embedding head. We investigate the effects of scale gates and embedding head dimensions, with and without L2 normalization, on model performance. As shown in Table 4, for Gemma-2B-it models, incorporating a scale gate generally enhances GPM performance across various embedding dimensions. L2 normalization on the embedding head output consistently improves models with scale gates. Interestingly, Gemma-2B-it-based models without L2 normalization or scale gates outperform those with L2 normalization but no scale gates. A plausible explanation for this phenomenon is that removing L2 normalization introduces additional degrees of freedom, particularly beneficial for models with smaller parameter spaces and high-dimensional embedding layers. This increased flexibility may allow the model to utilize its limited parametric capacity better, potentially leading to enhanced expressiveness and task-specific adaptability.
Table 4: Impact of the embedding head and the scale gate on GPM’s performance on RewardBench. Dim. represents the dimension of the embedding head. The highest average scores for each base model are in bold.
@@ E.1 Additional Ablation Studies
w. scale gate w. l2 2 77.37 73.46 85.00 85.50 80.33
w. scale gate w.o. l2 2 79.33 74.34 85.14 88.41 81.80
w. o. scale gate w. l2 2 78.49 71.27 85.68 86.13 80.39
w. o. scale gate w.o. l2 2 79.05 73.46 84.86 86.56 80.98
w. scale gate w. l2 4 78.77 72.59 85.44 84.82 80.43
w. scale gate w.o. l2 4 80.45 72.81 84.46 87.61 81.33
w. o. scale gate w. l2 4 79.61 70.39 85.00 86.84 80.46
w. o. scale gate w.o. l2 4 80.72 73.02 83.51 86.96 81.06
w. scale gate w. l2 6 79.61 75.66 85.27 88.61 82.29
w. scale gate w.o. l2 6 76.54 76.10 85.14 87.55 81.33
w. o. scale gate w. l2 6 79.61 71.05 85.81 87.74 81.05
w. o. scale gate w.o. l2 6 77.93 73.25 85.41 86.66 80.81
w. scale gate w. l2 8 78.49 74.34 84.19 86.95 81.00
w. scale gate w.o. l2 8 82.40 74.78 85.54 85.47 82.05
w. o. scale gate w. l2 8 77.09 72.15 86.08 85.41 80.18
w. o. scale gate w.o. l2 8 81.28 73.25 84.59 85.90 81.26
@@
w. scale gate w. l2 2 91.62 88.38 90.68 94.82 91.37
w. scale gate w.o. l2 2 93.85 86.84 90.68 91.60 90.74
w. o. scale gate w. l2 2 92.18 86.18 91.89 94.05 91.08
w. o. scale gate w.o. l2 2 93.30 87.94 91.22 93.55 91.50
w. scale gate w. l2 4 93.30 86.18 91.22 95.69 91.60
w. scale gate w.o. l2 4 94.13 86.18 89.86 90.55 90.18
w. o. scale gate w. l2 4 92.46 87.28 91.76 93.19 91.17
w. o. scale gate w.o. l2 4 93.58 86.40 90.95 95.33 91.56
w. scale gate w. l2 6 91.90 87.50 91.62 96.40 91.86
w. scale gate w.o. l2 6 93.02 85.75 91.08 91.31 90.29
w. o. scale gate w. l2 6 92.18 85.53 90.81 94.20 90.68
w. o. scale gate w.o. l2 6 93.30 87.94 90.95 90.90 90.77
w. scale gate w. l2 8 93.58 87.50 91.08 95.44 91.90
w. scale gate w.o. l2 8 93.02 87.06 90.81 92.20 90.77
w. o. scale gate w. l2 8 91.90 86.62 91.22 92.63 90.59
w. o. scale gate w.o. l2 8 93.02 87.72 90.68 90.16 90.39
E.2 Additional Experimental Results
More Results on Language Model Alignment. We further conduct additional evaluations of our fine-tuned models using various benchmarks. AlpacaEval 2.0 evaluation results are listed in Table 5, using GPT-4o-mini as evaluators. For MT-Bench, we used the default mode to let GPT-4 grade and give a score to the model’s answer, and the MT-Bench scores of aligned models are presented in Table 6.
Table 5: AlpacaEval 2.0 evaluation results. Base model: Llama3-8B-it, Evaluator: GPT-4o-mini. The results are grouped by the size and type of the RM or PM, and the number of iterations. Bold entries indicate that GPM outperforms BT RM under the same training settings.
Size Type Iter SPPO GPO
LC. WR WR Avg. Len LC. WR WR Avg. Len
base 23.07 32.26 1959 23.07 32.26 1959
2B BT RM 1 48.84 46.09 1939 53.15 49.94 1929
2 59.77 58.41 2032 66.19 64.88 2049
3 66.81 67.14 2136 71.75 71.68 2151
GPM 1 48.09 49.15 (+3.06) 2066 55.66 57.12 (+7.18) 2102
2 56.63 63.53 (+5.12) 2301 61.11 67.78 (+2.90) 2343
3 60.77 70.91 (+3.77) 2498 64.52 74.78 (+3.10) 2582
8B BT RM 1 45.24 36.95 1740 49.77 40.26 1702
2 56.24 50.36 1868 60.75 56.30 1933
3 63.71 58.38 1948 62.63 59.17 1969
GPM 1 46.84 41.42 (+4.47) 1861 53.12 46.64 (+6.38) 1850
2 58.03 56.07 (+5.71) 2029 59.86 60.37 (+4.07) 2115
3 61.64 63.42 (+5.04) 2385 62.51 67.48 (+8.31) 3249
Table 6: MT-Bench evaluation results. Base model: Llama3-8B-it, Evaluator: GPT-4. Bold entries indicate that GPM outperforms BT RM under the same training settings.
Size Type Iter SPPO GPO
1st 2nd Avg. 1st 2nd Avg.
base 8.31 7.77 8.03 8.31 7.77 8.03
2B BT RM 1 8.42 7.57 8.00 8.33 7.85 8.09
2 8.20 7.73 7.96 8.30 7.66 7.98
3 8.44 7.66 8.05 8.41 8.09 8.25
GPM 1 8.23 7.65 7.94 8.70 7.95 8.33
2 8.53 8.24 8.38 8.69 8.01 8.35
3 8.39 7.84 8.12 8.48 7.76 8.12
8B BT RM 1 8.44 8.10 8.27 8.41 7.85 8.13
2 8.75 7.85 8.30 8.73 7.83 8.28
3 8.34 7.99 8.17 8.68 7.83 8.26
GPM 1 8.43 7.94 8.18 8.29 7.90 8.10
2 8.51 8.05 8.28 8.26 7.99 8.13
3 8.47 7.76 8.12 7.57 7.51 7.54
Remark E.1.
We can have the following length-normalized (Meng et al., 2024) GPO (LN-GPO) learning objective (we report the empirical result in Table 7):
$\mathcal{L}_{\text{LN-GPO}}(\bm{\theta})=$ $\mathbb{E}_{\mathbf{x}\sim\mathcal{X},\mathbf{y}\sim\pi_{\bm{\theta}_{t}}(\cdot\mid\mathbf{x})}\Bigg{[}\Bigg{(}{\frac{1}{|y|}}\log\left(\frac{\pi_{\bm{\theta}}(\mathbf{y}\mid\mathbf{x})}{\pi_{\bm{\theta}_{t}}(\mathbf{y}\mid\mathbf{x})}\right)-\frac{1}{\beta}\left(\widehat{s}\left(\mathbf{y}\succ\pi_{\bm{\theta}_{t}}\mid\mathbf{x}\right)-\log Z_{\pi_{\bm{\theta}_{t}}}(\mathbf{x})\right)\Bigg{)}^{2}\Bigg{]}.$ (E.1)